9:00a – 10:50p
- Finish O/N E10 imaging
- Hao imaging
- set up O/N imaging of E12 (600 kb G) with 20:30,000 bead dilution.
- Qualitative observations: some G regions quite large, seem pretty decompact. Others more compact?
- Laser power dropped to 50 mW progammed power at some point randomly during movie _0_1? Reset to 500. How did this happen?! Good thing I caught it early.
- Transfered E08 and E10 data to ProBox 3 and deleted from STORM4-temp.
- working through G6 data
- almost all regions horribly out of focus. Calibration movies look good. Conventional images nice and clear in calibration. No dots evident except the feducial beads in main movies. This is probably all due to the horrible jumping of the stage in the z-scan. Need to slow down scan and scan smaller regions.
- Deleted out of focus G6 data.
- Running SplitDax on E10 data (should just fit in remaining space on ProBox3).
- working on talk for Rowland Center
- exploring model parameters. Trying to find regime where Ph + Ph-flag clusters are not also larger in the mutant due to concentration effects.
- Fixed a long-range bond bug due to computing
- Cluster size is limited by pool of free molecules, not by available binding sites.